# Embodied Arena: A Comprehensive, Unified, and Evolving Evaluation Platform for Embodied AI

### v1

@@ 3 Systematic Embodied Capability Taxonomy
Embodied Navigation RxR-CE (Yang et al., 2025c, Zhang et al., 2024c) Embodied Navigation: Instruction Navigation
Embodied Task Planning ET-Plan-Bench (Zhang et al., 2024d) Embodied Task Planning: Spatial Reference Planning, Temporal Reference Planning
@@ 4.1 Tasks, Benchmarks, and Data
Specifically, for Embodied Question Answering, we consider two types of benchmarks: 2D question answering and 3D question answering. The 2D question answering benchmarks include OpenEQA (Majumdar et al., 2024), VSI-Bench (Yang et al., 2024), and ERQA (Abeyruwan et al., 2025), etc., and the 3D question answering benchmarks include the representative ScanQA (Azuma et al., 2022), Scan2Cap (Chen et al., 2021), and SQA3D (Ma et al., 2023), etc. For Embodied Navigation, we select the classic object navigation benchmarks MP3D (Chang et al., 2017), HM3D (Ramakrishnan et al., 2021), EB-Navigation (Yang et al., 2025c), and the instruction navigation benchmarks R2R-CE (Krantz et al., 2020) and RxR-CE (Krantz et al., 2020, Zhang et al., 2024c). For Embodied Task Planning, we consider EB-ALFRED, EB-Habitat (Yang et al., 2025c), and ET-Plan-Bench (Zhang et al., 2024d), which are more diverse in task types.
@@ 4.2.1 General Multimodal Large Models
  • ByteDance: LongVA-7B (Zhang et al., 2024e), LLaVA-OneVision (Li et al., 2024a), LLaVA-NeXT-Video (Zhang et al., 2025c), pllava-7b (Xu et al., 2024a)
@@
  • Shanghai AI Lab: InternVL3 (Zhu et al., 2025),InterVL2.5 (Chen et al., 2024c), InternVL2, ChatScene (Zhang et al., 2024a)
@@ 4.2.2 Embodied AI Models
  • BAAI: Navid (Zhang et al., 2024c), UniNavid (Zhang et al., 2024b), RoboBrain1.0-7B (Ji et al., 2025), RoboBrain2.0-7B, RoboBrain2.0-32B (Team et al., 2025), MapNav (Zhang and Heidari, 2025)
  • Shanghai AI Lab: VeBrain (Luo et al., 2025), VLN-R1 (Qi et al., 2025b), navillm (Zheng et al., 2024), StreamVLN (Wei et al., 2025a)
@@
  • Peking University: UniNavid (Zhang et al., 2024b), Navid (Zhang et al., 2024c)
@@ 4.4.1 Evaluation Metrics
  • Task Completion Success Rate: Applied to benchmarks such as EB-ALFRED, EB-Habitat, and ET-Plan-Bench (Zhang et al., 2024d), measuring the percentage of successfully completed task sequences
@@ 6 Insights from Embodied Arena Evaluation and Leaderboards
    Core Finding: As illustrated in Figure 4, our evaluation reveals a significant phenomenon that massive general models achieve strong performance through large model size and broad knowledge, while specialized models excel through targeted embodied training. Large commercial general models, e.g., GPT-o3 (OpenAI, 2025), Gemini-2.5-Pro (Google DeepMind, 2024), Claude-3.7 (Anthropic, 2024), leverage their massive scale to achieve overall 10-20% performance advantages across most benchmarks, demonstrating that model size and extensive pre-training data clearly matter. However, when we compare models of similar scales, specialized embodied models consistently outperform the general multimodal models. RoboBrain2.0-32B (BAAI, 2025) achieves 73.59% on Where2Place versus GPT-o3’s 33.46%, while navigation specialists like StreamVLN (Wei et al., 2025a) and UniNavid (Zhang et al., 2024b) reach 30-57% success rates while the general models of similar sizes such as InternVL3 (Zhu et al., 2025) achieve less than 10% success rates.
@@
    Core Finding: Based on the comprehensive ranking of the models in embodied capabilities (e.g., object perception, spatial perception, temporal perception, embodied knowledge, and embodied reasoning) and downstream tasks (e.g., embodied navigation, embodied task planning), we find that models’ embodied capabilities are highly positively correlated with downstream task performance (Spearman’s rank correlation coefficient $\rho=0.80$, $p<0.0001$). Moreover, each embodied capability shows a significant positive correlation with downstream task performance, with correlation coefficients $\rho$ ranging from 0.73 to 0.83. Among these, embodied knowledge demonstrates the strongest positive correlation ($\rho=0.83$, $p<0.0001$), and the remaining capabilities show stable correlation coefficients around 0.75. All p-values for the above analyses are less than 0.0001, reaching an extremely high level of statistical significance. Notably, the strength of this correlation is significantly influenced by how the model is applied in downstream tasks: in an end-to-end manner, the model’s embodied capabilities and downstream task performance exhibit only a moderate correlation ($\rho=0.40$, $p>0.08$), which does not reach statistical significance; However, in a task-oriented agentic framework manner (i.e., integrating the general models into agentic frameworks targeted for downstream embodied tasks), the two exhibit a strong positive correlation ($\rho=0.79$, $p<0.0001$). In addition, we find that when general models are applied in an end-to-end manner, their overall success rate in navigation tasks is only 5.80%. In contrast, when using task-oriented agentic frameworks (such as EmbodiedBench (Yang et al., 2025c) and ET-Plan-Bench (Zhang et al., 2024d)), general models’ success rates in navigation and task planning tasks increase to 36.21% and 40.08%, respectively.
@@
    Core Finding: For the same model architecture, increasing the number of parameters can improve performance on specific embodied benchmarks. General multimodal models show more consistent evidence of this, whereas for embodied models, this phenomenon is inconsistent across models and capabilities. For instance, InternVL3-38B > InternVL3-14B > InternVL3-8B (Zhu et al., 2025) across all three fundamental capabilities. Similar trends are observed for RoboBrain2.0 (32B vs. 7B) (BAAI, 2025) and Qwen2.5-VL-Instruct (7B vs. 3B) (Bai et al., 2025) on embodied QA and embodied task planning. Nevertheless, the scaling effect is not consistent: in embodied navigation, smaller models (RoboBrain2.0-7B and Qwen2.5-VL-3B-Instruct) achieve better performance than their larger counterparts. Moreover, from the perspective of embodied data, increasing the amount of task-specific data can significantly improve models’ specific capabilities. For example, StreamVLN (Wei et al., 2025a), NaVILA (Zheng et al., 2024), and UniNavid (Zhang et al., 2024b) outperform GPT-o3 (OpenAI, 2025) in instruction-following navigation. However, the constructed embodied datasets usually fail to deliver consistent performance improvements in all capabilities. For example, Embodied-R1 (Yuan et al., 2025b) and SpaceR (Ouyang et al., 2025), trained on their respective embodied datasets, surpass the base model Qwen-2.5-VL-3B-Instruct in some capabilities. However, they also suffer from performance drops in others.
@@
    Core Finding: E2E frameworks show that VLN-specialized models demonstrate substantial performance advantages over general multimodal models through targeted architectural innovations and domain-specific training data. The top-performing VLN models — StreamVLN (Wei et al., 2025a) (54.90%), NaVILA (Zheng et al., 2024) (51.65%), and UniNavid (Zhang et al., 2024b) (47.85%) — achieve dramatically higher success rates compared to leading general models like Claude-3.7-Sonnet (Anthropic, 2024) (20.17%), Gemini-2.5-Pro (Google DeepMind, 2024) (27.37%), and GPT-4o (OpenAI, 2024) (21.67%). Notably, specialized VLN models dominate the entire top-5 rankings, with even mid-tier VLN models like MapNav (Zhang et al., 2025a) (36.15%) and Navid (Zhang et al., 2024c) (30.60%) outperforming most general foundation models. This performance gap is particularly pronounced in navigation-specific metrics, where StreamVLN (Wei et al., 2025a) achieves 56.90% on VLN-CE R2R compared to general models that struggle to exceed 25% success rates. And agentic frameworks demonstrate the potential to alleviate these built-in model limitations: specialized agentic frameworks like OmniEVA achieve top performance (59.10% on MP3D, 74.20% on HM3D) and OVRL (Yadav et al., 2023b) also achieve competitive results (62.00% on HM3D).
    In-depth Analysis: This performance gap between VLN-specialized and general models stems from architectural and training differences addressing embodied navigation challenges. VLN-specific architectures demonstrate superior performance by leveraging historical frames rather than relying solely on current frames, e.g., models like NaVid (Zhang et al., 2024c) and UniNavid (Zhang et al., 2024b) allocate more tokens to current frames for improved decision accuracy. The combination of RxR and R2R datasets with Habitat simulation enables large-scale vision-language-action data construction for VLN, allowing effective supervised fine-tuning and strong validation performance on unseen splits. Building upon these capabilities, designing efficient hierarchical agentic frameworks represents a promising direction for leveraging VLN capabilities. Such frameworks could decompose complex navigation tasks into subtasks, integrate multi-modal reasoning with spatial planning, and provide error recovery mechanisms. Agentic approaches activate and amplify existing foundation model capabilities through external reasoning pipelines rather than requiring costly model retraining or architectural modifications. This enables consistent performance enhancement through pipeline optimization, where new knowledge sources, memory architectures, and reasoning strategies can be systematically integrated without modifying the foundation model itself.

### v2

@@ 3 Systematic Embodied Capability Taxonomy
Embodied Navigation RxR-CE (Yang et al., 2025c, Zhang et al., 2024b) Embodied Navigation: Instruction Navigation
Embodied Task Planning ET-Plan-Bench (Zhang et al., 2024c) Embodied Task Planning: Spatial Reference Planning, Temporal Reference Planning
@@ 4.1 Tasks, Benchmarks, and Data
Specifically, for Embodied Question Answering, we consider two types of benchmarks: 2D question answering and 3D question answering. The 2D question answering benchmarks include OpenEQA (Majumdar et al., 2024), VSI-Bench (Yang et al., 2024), and ERQA (Abeyruwan et al., 2025), etc., and the 3D question answering benchmarks include the representative ScanQA (Azuma et al., 2022), Scan2Cap (Chen et al., 2021), and SQA3D (Ma et al., 2023), etc. For Embodied Navigation, we select the classic object navigation benchmarks MP3D (Chang et al., 2017), HM3D (Ramakrishnan et al., 2021), EB-Navigation (Yang et al., 2025c), and the instruction navigation benchmarks R2R-CE (Krantz et al., 2020) and RxR-CE (Krantz et al., 2020, Zhang et al., 2024b). For Embodied Task Planning, we consider EB-ALFRED, EB-Habitat (Yang et al., 2025c), and ET-Plan-Bench (Zhang et al., 2024c), which are more diverse in task types.
@@ 4.2.1 General Multimodal Large Models
  • ByteDance: LongVA-7B (Zhang et al., 2024d), LLaVA-OneVision (Li et al., 2024a), LLaVA-NeXT-Video (Zhang et al., 2025c), pllava-7b (Xu et al., 2024a)
@@
  • Shanghai AI Lab: InternVL3 (Zhu et al., 2025),InterVL2.5 (Chen et al., 2024c), InternVL2,
@@ 4.2.2 Embodied AI Models
  • BAAI: Navid (Zhang et al., 2024b), UniNavid (Zhang et al., 2024a), RoboBrain1.0-7B (Ji et al., 2025), RoboBrain2.0-7B, RoboBrain2.0-32B (Team et al., 2025), MapNav (Zhang and Heidari, 2025)
  • Shanghai AI Lab: VeBrain (Luo et al., 2025), VLN-R1 (Qi et al., 2025b), StreamVLN (Wei et al., 2025a)
@@
  • Peking University: UniNavid (Zhang et al., 2024a), Navid (Zhang et al., 2024b)
@@ 4.4.1 Evaluation Metrics
  • Task Completion Success Rate: Applied to benchmarks such as EB-ALFRED, EB-Habitat, and ET-Plan-Bench (Zhang et al., 2024c), measuring the percentage of successfully completed task sequences
@@ 6 Insights from Embodied Arena Evaluation and Leaderboards
    Core Finding: As illustrated in Figure 4, our evaluation reveals a significant phenomenon that massive general models achieve strong performance through large model size and broad knowledge, while specialized models excel through targeted embodied training. Large commercial general models, e.g., GPT-o3 (OpenAI, 2025), Gemini-2.5-Pro (Google DeepMind, 2024), Claude-3.7 (Anthropic, 2024), leverage their massive scale to achieve overall 10-20% performance advantages across most benchmarks, demonstrating that model size and extensive pre-training data clearly matter. However, when we compare models of similar scales, specialized embodied models consistently outperform the general multimodal models. RoboBrain2.0-32B (BAAI, 2025) achieves 73.59% on Where2Place versus GPT-o3’s 33.46%, while navigation specialists like StreamVLN (Wei et al., 2025a) and UniNavid (Zhang et al., 2024a) reach 30-57% success rates while the general models of similar sizes such as InternVL3 (Zhu et al., 2025) achieve less than 10% success rates.
@@
    Core Finding: Based on the comprehensive ranking of the models in embodied capabilities (e.g., object perception, spatial perception, temporal perception, embodied knowledge, and embodied reasoning) and downstream tasks (e.g., embodied navigation, embodied task planning), we find that models’ embodied capabilities are highly positively correlated with downstream task performance (Spearman’s rank correlation coefficient $\rho=0.80$, $p<0.0001$). Moreover, each embodied capability shows a significant positive correlation with downstream task performance, with correlation coefficients $\rho$ ranging from 0.73 to 0.83. Among these, embodied knowledge demonstrates the strongest positive correlation ($\rho=0.83$, $p<0.0001$), and the remaining capabilities show stable correlation coefficients around 0.75. All p-values for the above analyses are less than 0.0001, reaching an extremely high level of statistical significance. Notably, the strength of this correlation is significantly influenced by how the model is applied in downstream tasks: in an end-to-end manner, the model’s embodied capabilities and downstream task performance exhibit only a moderate correlation ($\rho=0.40$, $p>0.08$), which does not reach statistical significance; However, in a task-oriented agentic framework manner (i.e., integrating the general models into agentic frameworks targeted for downstream embodied tasks), the two exhibit a strong positive correlation ($\rho=0.79$, $p<0.0001$). In addition, we find that when general models are applied in an end-to-end manner, their overall success rate in navigation tasks is only 5.80%. In contrast, when using task-oriented agentic frameworks (such as EmbodiedBench (Yang et al., 2025c) and ET-Plan-Bench (Zhang et al., 2024c)), general models’ success rates in navigation and task planning tasks increase to 36.21% and 40.08%, respectively.
@@
    Core Finding: For the same model architecture, increasing the number of parameters can improve performance on specific embodied benchmarks. General multimodal models show more consistent evidence of this, whereas for embodied models, this phenomenon is inconsistent across models and capabilities. For instance, InternVL3-38B > InternVL3-14B > InternVL3-8B (Zhu et al., 2025) across all three fundamental capabilities. Similar trends are observed for RoboBrain2.0 (32B vs. 7B) (BAAI, 2025) and Qwen2.5-VL-Instruct (7B vs. 3B) (Bai et al., 2025) on embodied QA and embodied task planning. Nevertheless, the scaling effect is not consistent: in embodied navigation, smaller models (RoboBrain2.0-7B and Qwen2.5-VL-3B-Instruct) achieve better performance than their larger counterparts. Moreover, from the perspective of embodied data, increasing the amount of task-specific data can significantly improve models’ specific capabilities. For example, StreamVLN (Wei et al., 2025a), NaVILA (Zheng et al., 2024), and UniNavid (Zhang et al., 2024a) outperform GPT-o3 (OpenAI, 2025) in instruction-following navigation. However, the constructed embodied datasets usually fail to deliver consistent performance improvements in all capabilities. For example, Embodied-R1 (Yuan et al., 2025b) and SpaceR (Ouyang et al., 2025), trained on their respective embodied datasets, surpass the base model Qwen-2.5-VL-3B-Instruct in some capabilities. However, they also suffer from performance drops in others.
@@
    Core Finding: E2E frameworks show that VLN-specialized models demonstrate substantial performance advantages over general multimodal models through targeted architectural innovations and domain-specific training data. The top-performing VLN models — StreamVLN (Wei et al., 2025a) (54.90%), NaVILA (Zheng et al., 2024) (51.65%), and UniNavid (Zhang et al., 2024a) (47.85%) — achieve dramatically higher success rates compared to leading general models like Claude-3.7-Sonnet (Anthropic, 2024) (20.17%), Gemini-2.5-Pro (Google DeepMind, 2024) (27.37%), and GPT-4o (OpenAI, 2024) (21.67%). Notably, specialized VLN models dominate the entire top-5 rankings, with even mid-tier VLN models like MapNav (Zhang et al., 2025a) (36.15%) and Navid (Zhang et al., 2024b) (30.60%) outperforming most general foundation models. This performance gap is particularly pronounced in navigation-specific metrics, where StreamVLN (Wei et al., 2025a) achieves 56.90% on VLN-CE R2R compared to general models that struggle to exceed 25% success rates. And agentic frameworks demonstrate the potential to alleviate these built-in model limitations: specialized agentic frameworks like OmniEVA achieve top performance (59.10% on MP3D, 74.20% on HM3D) and OVRL (Yadav et al., 2023b) also achieve competitive results (62.00% on HM3D).
    In-depth Analysis: This performance gap between VLN-specialized and general models stems from architectural and training differences addressing embodied navigation challenges. VLN-specific architectures demonstrate superior performance by leveraging historical frames rather than relying solely on current frames, e.g., models like NaVid (Zhang et al., 2024b) and UniNavid (Zhang et al., 2024a) allocate more tokens to current frames for improved decision accuracy. The combination of RxR and R2R datasets with Habitat simulation enables large-scale vision-language-action data construction for VLN, allowing effective supervised fine-tuning and strong validation performance on unseen splits. Building upon these capabilities, designing efficient hierarchical agentic frameworks represents a promising direction for leveraging VLN capabilities. Such frameworks could decompose complex navigation tasks into subtasks, integrate multi-modal reasoning with spatial planning, and provide error recovery mechanisms. Agentic approaches activate and amplify existing foundation model capabilities through external reasoning pipelines rather than requiring costly model retraining or architectural modifications. This enables consistent performance enhancement through pipeline optimization, where new knowledge sources, memory architectures, and reasoning strategies can be systematically integrated without modifying the foundation model itself.